OpenStudy (khally92):

Anybody please Use the method of Lagrange multipliers to find the max and min values of g(s,t)= t^2 e^s given that s^2+t^2=3 . How are we assured that these extreme values exist? I am having Difficulty solving the system of lagrange equation Thank you.

OpenStudy (khally92):

2λs=t^2e^s 2λt=2te^s s^2+t^2=3

OpenStudy (khally92):

Thanks alot.

OpenStudy (irishboy123):

just ripping mechanically through the problem gives you something like this \(\large g(s,t)= e^s t^2, \qquad h(s,t) = s^2+t^2=3\) \(\large \nabla g = \lambda \nabla h \) [the central idea] \(\large \implies <e^s t^2, e^s 2t> = \lambda <2s, 2t>\) \(\large \implies \dfrac{e^st^2}{2s} = \dfrac{e^s2t}{2t}\) \(\large \implies 2s = t^2\) and from \(\large h(s,t) = s^2+t^2=3\) you have \(\large s = -3, 1\) and \(t = \dots\)

OpenStudy (irishboy123):

if you are interested, we can work through what the problem actually explores, in termsof the geometry/physical interpretation.

OpenStudy (khally92):

Sorry how did you get s=-3 isnt suppose to be 3

OpenStudy (khally92):

s^2+2s=3 s(s+2)=3 s=3 s=-2

OpenStudy (irishboy123):

\(s^2 + 2s = 3\) \(s^2 + 2s - 3 = 0\) \((s+ 3)(s-1) = 0\) \(s = -3, 1\)

OpenStudy (khally92):

And no t value for s=-3

OpenStudy (irishboy123):

no. according to me. can you see this in your head? the underlying function is being constrained by/ looked at around a circle or radius \(\sqrt{3}\)

OpenStudy (khally92):

Thank you.

OpenStudy (khally92):

Ok, can you please work me through the questions in terms of the geometry/physical interpretation.